Deneen bought 120 floor tiles and 3 cans of paint for $195. If a can of paint cost $3.50 more than each tile, find the cost of two cans of paint.
Answer by MathTherapy(10858)

Deneen bought 120 floor tiles and 3 cans of paint for $195. If a can of paint cost $3.50 more than each tile, find the cost of two cans of paint.

Cost of 1 tile = t Therefore, cost of a can of paint = t + 3.5

120t + 3(t + 3.5) = 195
120t + 3t + 10.5 = 195
123t = 184.5
t = $1.50

Since a can of paint costs $3.50 more than 1 tile, then a can of paint costs $1.50 + $3.50 = $5. Therefore, 2 cans of paint cost $10.